The length of a rectangular garden is 8 feet longer than its width. If the garden's perimeter is 204 feet, what is the area of the garden in square feet?
Answer by addingup(3677)

2L + 2W = 204 (1)
L = W + 8
substitute for L in (1)
2(W+8) + 2W = 204
Distribute the 2 (Distributive Properties of Multiplication)
2W+16+2W = 204
4W = 188
W = 47
~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2(47+8)+2(47) = 204
110+94 = 204 Correct